If Hubble's constant, H0, was measured to be 6.25 x 10-11 (years)-1, what would be the age of the universe?
Answer #1

Age of universe = 1/H

H = Hubble constant

Age of universe = 1/6.25E-11/years

= 0.16E+11 years = 16E+9year

= 16 billion years